If we know a quadratic function f has a zero at x = −1
and vertex at (1, 4), do we have enough information to
find a formula for this function? If your answer is yes,
find it; if not, give your reasons.

Answer by lwsshak3(11628)

If we know a quadratic function f has a zero at x = −1
and vertex at (1, 4), do we have enough information to
find a formula for this function? If your answer is yes,
find it; if not, give your reasons.
***
function is a parabola that opens down
Its basic form of equation:
(x-h)^2=-4p(y-k)
Plug in coordinates of vertex(1,4) and x-intercept(-1,0)
(-1-1)^2=-4p(0-4)
4=16p
p=4/16=1/4
4p=1
equation:(x-1)^2=-(y-4)
..
do we have enough information to
find a formula for this function? yes, as illustrated above